OpenStudy (anonymous):

Solve for x: ln (x^2+1)-3 ln x=ln(2)

OpenStudy (anonymous):

A ln B=ln(B^A); ln A+ ln B= ln AB; ln A- ln B= ln (A/B); ln A=lnB--->A=B